OpenStudy (anonymous):

I have to graph this polynomial f(x)=x^3-4x+1 Where do I start?

OpenStudy (hunus):

Do you know what x^3 looks like?

OpenStudy (anonymous):

yes, like if I put in a 1, x is still 1. If I put in a 2, x is 8

OpenStudy (hunus):

Let's start with x=0

OpenStudy (hunus):

\[f(0)=0^3 - 4*0 + 1 = 1\]

OpenStudy (hunus):

You're having to actually graph this, right?

OpenStudy (anonymous):

Yes

OpenStudy (hunus):

Do you know to what x-values are you having to graph it?

OpenStudy (anonymous):

So, with x=0, y just =-3, right?

OpenStudy (hunus):

With x=0, y=1

OpenStudy (hunus):

because zero cubed is zero and four times zero is still zero.

OpenStudy (anonymous):

Oh my goodness :/ I was adding it and driving myself nuts, I'm good, thanks:)

OpenStudy (anonymous):

I've been doing toooo much homework sorry

OpenStudy (hunus):

You're fine lol

OpenStudy (hunus):

You still need help with this one?

OpenStudy (anonymous):

No, I can get my five points now and make my graph. Thanks.

OpenStudy (anonymous):

(0, 1) (1,3)(-1,4)(2,1)(-2,1)? no

OpenStudy (hunus):

For x = 1 1^3 - 4*1 + 1 = 1 - 4 + 1 = -2

OpenStudy (hunus):

(0, 1) (1, -2)

OpenStudy (anonymous):

Thanks

OpenStudy (hunus):

Yup